Mia is 4 years older than her sister Taylor. Mia’s brother is nine years younger than their sister Taylor. The sum of their ages is 64 years.
Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

Write an equation for each statement
:
Mia is 4 years older than her sister Taylor.
m = t + 4
Mia’s brother is nine years younger than their sister Taylor.
b = t - 9
The sum of their ages is 64 years.
m + t + b = 64
replace m with (t+4) and replace b with (t-9)
(t+4) + t + (t-9) = 64
3t - 5 = 64
3t = 64 + 5
3t = 69
t = 69/3
t = 23 yrs is Taylor's name
then
m = 23 + 4
m = 27 yrs is Mia's age
and
b = 23 - 9
b = 14 yrs is brother's age
:
:
Check by find the sum of their ages
23 + 27 + 14 = 74

Answer by ikleyn(52787)

.

This post does not deserve any consideration, since it came with no question.